ABC Corp. has an ROE of 5% and reinvests 40% of its net income. ABC has just paid an annual dividend of $0.29.

ABC stock has a beta of 1.3. The risk-free rate is 3.6% and the expected return on the market portfolio is 7%.

1) What is the appropriate discount rate?

2) What is the expected growth rate of dividends?

3) What is the intrinsic value (fair price) of ABC stock?
Answer #1

Answer 1)

Appropriate Rate is given by

CAPM = Rf + β(rm- Rf)

Where Rf is the risk free rate of return
Rm is the market Return

= 3.6% + 1.3(7%-3.6%)

= 8.02%

Answer 2 )

Growth Rate of Dividend = Retention Ratio * RoE

= 40% * 5%

= 2%

Answer 3 )

Intrinsic Value = \frac{D_{0}(1+g)}{Rate of Return- g }

= \frac{0.29(1+0.02)}{0.0802- 0.02 }

= 0.2958/0.0602

= $ 4.9136

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.